# Taming Anomalies with Down-Up Sampling Networks: Group Center Preserving Reconstruction for 3D Anomaly Detection

### v1

@@ Abstract.
This work was done by Hanzhe Liang at Shenzhen University, please email Hanzhe for any questions.
submissionid: 4conference: Make sure to enter the correct conference title from your rights confirmation email; 2025; Dublin, Irelandccs: Computing methodologies Visual inspection
@@ 3.1. Patch Augmentation with Noise Generation
Given a point cloud $\mathcal{P}$in $\mathbb{R}^{N\times 3}$, we partition it into $G$ patches $\{P_{i}\}_{i=1}^{G}$ via farthest point sampling (FPS) and K-nearest neighbors (KNN). Each patch $P_{i}$in $\mathbb{R}^{(K+1)\times 3}$ is composed by $P_{i}$ = $\{c_{i}\cup\mathcal{N}_{K}(c_{i})\}$, where $c_{i}$ denotes the center point selected by the FPS, and $\mathcal{N}_{K}(c_{i})$ means the $K$ nearest neighbors of $c_{i}$. To diversify the point cloud, Gaussian noises are injected into each group with two controllable noise parameters $\alpha$ and $\beta$. The patch after noise injection can be formulated as
@@
=$=\{GSN(c_{i},\alpha)\cup GSN(\mathcal{N}_{K}(c_{i}),\beta)\},$
=$=\{\tilde{c}_{i}\cup\tilde{\mathcal{N}}_{K}(c_{i})\},$
@@ 3.2. Center-preserved Down-sampling with Down-Net
[…]
@@ 3.3. Multiscale Reconstruction with Up-Net
where $p_{i}^{l}$ denotes a point in the $l$-th scale down-sampled point cloud $\tilde{\mathcal{P}}^{l}$, $p_{j}^{l-1}$ and $F_{j}^{l-1}$ represents the points from the neighbor set $\mathcal{N}^{l-1}(p_{i}^{l})$ of $p_{i}^{l}$ in the $l-1$-th scale and the corresponding extracted feature, respectively, and the symbols $MaxPool$ and $MLP$ stand for the operator of max-pooling and the feature extractor MLP, respectively.
@@ 4.4. Parameter Sensitivity Analysis
4.5. Robustness and Stability Analysis
Equipment instability and improper operations may result in low-quality point clouds. We considered two realistic challenges: incomplete point clouds and noisy point clouds.
Table 9. Results of robustness experiments.
Scale 1/2 1/4 1/6 1/8 Origin
O-AUROC 0.788 0.793 0.772 0.758 0.797
P-AUROC 0.710 0.701 0.654 0.640 0.712
O-AUPR 0.798 0.800 0.764 0.735 0.802
P-AUPR 0.238 0.232 0.210 0.183 0.244
(a) Incomplete point clouds.
Deviations 0.001 0.003 0.005 0.007 0.009 Origin
O-AUROC 0.796 0.780 0.769 0.751 0.713 0.797
P-AUROC 0.711 0.705 0.699 0.696 0.674 0.712
O-AUPR 0.787 0.782 0.778 0.759 0.720 0.802
P-AUPR 0.237 0.231 0.224 0.218 0.184 0.244
(b) Noisy point clouds.
Incomplete Point Clouds. An incomplete point cloud means that some points are missed during data collection, and we examined three scales: point cloud with a down-sampling rate of $1/2$, $1/4$, $1/6$, and $1/8$. As shown in Table 9, our method maintained outstanding performance when facing incomplete point clouds. For instance, the anomaly detection performance only decreased by 3.9%, 7.2%, 6.7%, and 6.1% in a down-sampling rate of $1/8$, which is still higher than the second-best comparative methods in O-AUROC, demonstrating the robustness of our method in the conditions of incomplete point cloud.
Noisy Point Clouds. We simulated noisy point clouds by injecting Gaussian noise with a standard deviation of 0.001, 0.003, 0.005, 0.007, or 0.009, respectively. As seen in Table 9, our approach achieved good performance in noisy situations, For example, when the standard deviation was 0.009, our approach still maintained a good performance with O-AUROC, P-AUROC, O-AUPR, and P-AUPR of 71.3%, 67.4%, 72.0% and 18.4%, respectively, which are still competitive with some other methods. This may be due to that our method employs a group center-based reconstruction mechanism and that our Down-Net has the ability to remove noise from points or patches.
@@ 5. Conclusion
Acknowledge
This work was supported by National Undergraduate Training Program for Innovation and Entrepreneurship (S202510590118).

### v2

@@ Abstract.
Corresponding author: Can Gao. Email: davidgao@szu.edu.cn
submissionid: 4conference: Make sure to enter the correct conference title from your rights confirmation email; 2025; Dublin, Irelandccs: Computing methodologies Visual inspectionjournalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3754492isbn: 979-8-4007-2035-2/2025/10
@@ 3.1. Patch Augmentation with Noise Generation
Given a point cloud $\mathcal{P}$ $\in$ $\mathbb{R}^{N\times 3}$, we partition it into $G$ patches $\{P_{i}\}_{i=1}^{G}$ via farthest point sampling (FPS) and K-nearest neighbors (KNN). Each patch $P_{i}$ $\in$ $\mathbb{R}^{(K+1)\times 3}$ is composed by $P_{i}$ = $\{c_{i}\cup\mathcal{N}_{K}(c_{i})\}$, where $c_{i}$ denotes the center point selected by the FPS, and $\mathcal{N}_{K}(c_{i})$ means the $K$ nearest neighbors of $c_{i}$. To diversify the point cloud, Gaussian noises are injected into each group with two controllable noise parameters $\alpha$ and $\beta$. The patch after noise injection can be formulated as
@@
$=\{GSN(c_{i},\alpha)\cup GSN(\mathcal{N}_{K}(c_{i}),\beta)\},$
$=\{\tilde{c}_{i}\cup\tilde{\mathcal{N}}_{K}(c_{i})\},$
@@ 3.2. Center-preserved Down-sampling with Down-Net
where $\mathcal{P}_{c}$ and $\tilde{\mathcal{P}}_{c}$ denote the original and predicted group center set, respectively, and $\parallel\cdot\parallel$ represents the 2-norm between two points.
@@ 3.3. Multiscale Reconstruction with Up-Net
where $p_{i}^{l}$ denotes a point in the $l$-th scale down-sampled point cloud $\tilde{\mathcal{P}}^{l}$, $p_{j}^{l-1}$ and $F_{j}^{l-1}$ represents the points from the neighbor set $\mathcal{N}^{l-1}(p_{i}^{l})$ of $p_{i}^{l}$ in the l1l-1-th scale and the corresponding extracted feature, respectively, and the symbols $MaxPool$ and $MLP$ stand for the operator of max-pooling and the feature extractor MLP, respectively.
@@ 5. Conclusion
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (Grant Nos. 62476171 and 62206122), Guangdong Basic and Applied Basic Research Foundation (Grant No. 2024A1515011367), National Undergraduate Training Program for Innovation and Entrepreneurship (Grant No. S202510590118), Guangdong Provincial Key Laboratory (Grant No. 2023B1212060076), Tencent “Rhinoceros Birds” - Scientific Research Foundation for Young Teachers of Shenzhen University, and the Shenzhen Institute of Artificial Intelligence and Robotics for Society.